A battery is connected to three resistors R1=8.0, R2=12 and R3=5.0 . Find the equivalent resistance in the circuit
1
Expert's answer
2019-11-21T10:01:57-0500

If the resistors are connected in series:


"R=R_1+R_2+R_3=8.0+12+5.5=25.5\space\Omega."

If they are all connected in parallel:


"1/R=1/R_1+1/R_2+1/R_3,\\
\space\\
R=\frac{R_1R_2R_3}{R_1R_2+R_2R_3+R_1R_3}=2.45\space\Omega."